(7+ (square root of 3)(7 - square root of 3) I didn't know how to type the square root sign
Answer by uma(370)


(7 + sqrt 3) ( 7 - sqrt3)
= 7^2 - (sqrt3)^2 [as (a + b)(a-b) = a^2 - b^2 ]
= 49 - 3
= 46

Good Luck